ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 179, 2010

Adachi, J. Global stability of special multi-flags  29

Adin, R. M., A. Postnikov and Y. Roichman A Gelfand model for wreath products

381

Antić, M. and L. Vrancken Sequences of minimal surfaces in S^{2n+1}  493

Assani, I. and Z. Buczolich The (L^{p}, L^{q}) bilinear Hardy-Littlewood function for the tail

173

Berkovich, Y. On finite p-groups with few nonabelian subgroups of order p^{p} and exponent p

189

Brandão, A. P., Jr., P. Koshlukov, A. Krasilnikov and É. A. da Silva The central polynomials for the Grassmann algebra

127

Buczolich, Z. See I. Assani Burdges, J. and A. Deloro Weyl groups of small groups of finite

Morley rank

403

Chow, C.-O. and T. Mansour Counting derangements, involutions and unimodal elements in the wreath product C_{r} ≀ 𝔖_{n}

425

Cilleruelo, J. and T. H. Lê On a question of Sárközy on gaps of product sequences

285

Coudene, Y. and B. Schapira Generic measures for hyperbolic flows on non-compact spaces

157

Deloro, A. See J. Burdges Fonf, V. P., M. Lin and P. Wojtaszczyk A non-reflexive Banach

space with all contractions mean ergodic

479

Fried, M. D. Alternating groups and moduli space lifting invariants  57

Grünbaum, B., T. Kaiser, D. Král’ and M. Rosenfeld Equipartite polytopes

235

Guo, X. Y. See Q. J. Kong Hasson, A., A. Onshuus and Y. Peterzil Definable one dimen-

sional structures in o-minimal theories

297

Hasson, A., A. Onshuus and Y. Peterzil Definable structures in o-minimal theories: one dimensional types

363

Kaiser, T. See B. Grünbaum Kong, Q. J. and X. Y. Guo On an extension of a theorem on con-

jugacy class sizes

279

Koshlukov, P. See A. P. Brandão, Jr.

Král’, D. See B. Grünbaum Krasilnikov, A. See A. P. Brandão, Jr. Kriete, T. L., B. D. MacCluer and J. L. Moorhouse Compo-

sition operators within singly generated composition C^{∗}-algebras

449

Lê, T. H. See J. Cilleruelo Lazar, A. J. Quotient spaces determined by algebras of continuous

functions

145

Lin, M. See V. P. Fonf MacCluer, B. D. See T. L. Kriete Mansour, T. See C.-O. Chow Moorhouse, J. L. See T. L. Kriete Onshuus, A. See A. Hasson Peterzil, Y. See A. Hasson Postnikov, A. See R. M. Adin Roichman, Y. See R. M. Adin Rosenfeld, M. See B. Grünbaum

Sanders, T. Chowla’s cosine problem  1

Schapira, B. See Y. Coudene da Silva É. A. See A. P. Brandão, Jr. Timoshin, S. A. Wiener criterion on metric spaces: Boundary regu-

larity in axiomatic and Poincaré-Sobolev spaces

211

Vrancken, L. See M. Antić Wojtaszczyk, P. See V. P. Fonf

Wolf, J. The structure of popular difference sets  253